# Gesplat: Robust Pose-Free 3D Reconstruction via Geometry-Guided Gaussian Splatting

### v1

@@ III-D Flow-based Depth Regularization
where $I_{i}$ and $I_{j}$ are the matching image pair. Given a point $p_{i}$ in image $I_{i}$, the predicted corresponding point $\widehat{p}_{j}$ in image $I_{j}$ is estimated as:
@@
Due to estimation errors, $\widehat{p}_{j}$ may not accurately lie on the epipolar line $l_{p_{i}}$. Therefore, we select the perpendicular foot $\overline{p}_{j}$ as the approximate estimation of $\widehat{p}_{j}$ :
@@
where $(\widehat{x}_{j},\widehat{y}_{j})$ are the coordinates of point $\widehat{p}_{j}$, and $\{a,b,c\}$ are the parameters of the epipolar line $l_{p_{i}}$. Then the depth value for point $p_{i}$ is calculated as:
@@
Cross-view Depth Blending. Given image set $\{I_{i}\mid i=1,2,\cdots,N\}$, for a point $p_{i}$ in $I_{i}$, we compute the predicted matching points $\{\overline{p}_{j}\mid j\neq i\}$ and corresponding depths in other N1N-1 views. Although $\overline{p}_{j}$ lies on the epipolar line, the distance error that $\overline{p}_{j}$ deviate from the true correspondence $\widehat{p}_{j}$ still persist. To handle the variance of depth estimation across views, we introduce a cross-view depth blending strategy.
Depth is represented as the distance information from a pixel point to the camera. Given $O_{i}$ as the camera center of view $i$, we define a reference distance $dis_{ref}$ as the distance from the 3D point $\overline{P}_{j}$ to $O_{i}$. And the projection distance $dis_{pro}$ in view $j$ is the distance from the projected point $\overline{p}_{j}$ to the epipole $o_{ij}$. The distance between $\overline{p}_{j}$ and the true matching point $p_{j}$ is denoted as $\Delta dis_{pro}$, corresponding to a 3D distance $\Delta dis_{ref}$. A smaller $\Delta dis_{ref}$ indicates a more accurate depth estimation. Through the fundamental theorem of calculus, the transformation between $\Delta dis_{ref}$ and $\Delta dis_{pro}$ is computed as:
@@
where $dis_{ref}^{\prime}(dis_{pro})$ is the gradient of $dis_{ref}(dis_{pro})$, indicating the sensitivity of depth to changes of $dis_{pro}$. A smaller $dis_{ref}^{\prime}(dis_{pro})$ implies higher confidence in the estimated depth. Using geometric relations, the formulation of $dis_{ref}^{\prime}(dis_{pro})$ is expressed as:
$dis_{ref}^{\prime}(dis_{pro})=\frac{\tan\beta\sin^{2}(\alpha+\theta)}{\min\theta\sin^{2}(\alpha+\beta)},$ (23)
@@
Depth Regularization. Given the rendered depth map $D$ and the flow-estimated depth map $\widehat{D}$, the depth loss is computed using loss function L1 :
@@
III-E Joint Optimization
To jointly optimize noisy point clouds and inaccurate camera poses, we introduce joint optimization, which updates both Gaussian and camera parameters via image loss. Given a Gaussian set $G$ and camera parameters $T$, we minimize the difference between ground truth and rendered images using gradient descent:
@@ III-E Joint Optimization
During testing, since predicted camera poses for novel views may still be noisy, we use the trained Gaussian model to refine test camera poses via Eq. (25) for 500 iterations per-image for more accurate evaluation.
III-F Training Objectives
Loss Function. The total training loss is formulated as :
@@ III-F Training Objectives
Training Details. We set $\lambda_{gp}=1.0$, $\lambda_{depth}=0$, $\lambda_{rg}=0$ for the first 2000 iterations, and then adjust to $\lambda_{depth}=0.1$, $\lambda_{rg}=0.3$ to avoid sub-optimization during early stage of training . Our model is built upon InstantSplat [19] and trained for 5000 iterations. The learning rate for the learnable depth $z$ starts at 0.1 and decays to $1.6\times 10^{-6}$. To mitigate initial errors and computational cost from the graph network, we only use it for Gaussian optimization in the final 200 iterations.
Figure 3: Qualititative comparision between Gesplatter and various pose-free methods on LLFF datasets(6 views).The reconstruction of our method is more accurate and exhibits finer details compared with other competitors.
TABLE I: Quantitative comparison on LLFF  (3,6,9 views). The best, second best, and third best entries are marked in red, orange, and yellow, respectively.
@@
IV Experiments
IV-A Experimental Settings
TABLE II: Quantitative comparison on Tank and Temples. All models are trained with six input views.The best result is highlighted in bold.
@@ IV-A Experimental Settings
Figure 4: Qualititative comparision on TNT(3,6,9 views).The reconstruction of our method is more accurate and exhibits finer details.
Datasets. Under sparse-view settings, we evaluate Gesplat using LLFF and Tanks and Temples datasets. LLFF provides a collection of real-world scenes with forward-facing camera trajectories, while Tanks and Temples is a large-scale dataset containing both indoor and outdoor scenes with complex environments. Both datasets consist of eight scenes each. Following InstantSplat, we randomly sample 24 images from each scene, with 12 images held out as test views (including the first and last images). From the remaining 12 images, we select 3 to 9 images for training under sparse-view conditions.
Baselines &\& Metrics. We compare Gesplat with pose-free approaches, including Nope-NeRF [17], CF-3DGS [18], and InstantSplat [19]. Among them, Nope-NeRF and CF-3DGS leverage monocular depth estimation and ground-truth camera intrinsics, while InstantSplat relies on DUSt3R to generate dense point clouds. All of the baselines are trained using official code with default settings. We evaluate the performance on novel view synthesis task using common metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [67], and Learned Perceptual Image Patch Similarity (LPIPS) [72].
Implementation Details. Our implementation is built on the PyTorch framework, and all experiments are conducted on one NVIDIA RTX 3090 GPU. For Nope-NeRF and CF-3DGS, we selete input images at the same resolution as original datasets. Using DUSt3R [20], InstantSplat predicts multi-view stereo depth maps at a resolution of 512. We train with a resolution of 518 × 392 using VGGT [21].
IV-B Experimental Results and Analysis
Results on LLFF. Quantitative comparisons are summarized in Tab. I Gesplat achieves the best performance across PSNR, SSIM, and LPIPS compared to other pose-free methods. Nope-NeRF utilizes Multilayer Perceptions (MLPs) to prioritize learning the global scene structure, which effectively suppresses floating artifacts but may produces images suffering from over-blur and requires a long training time. CF-3DGS builds upon 3DGS with local and global optimization, enabling more accurate geometric reconstruction and real-time rendering, yet its complex optimization process may lead to loss of fine details. As a feed-forward model, InstantSplat uses DUSt3R to estimate dense point clouds and camera poses, mitigating the information scarcity of 3DGS in sparse views and improving geometric consistency. However, it may occasionally generate inexistent details, reducing the reconstruction quality. By leveraging VGGT for dense 3D geometry recovery, our method uses hybrid Gaussian representation to learn scene geometry and refines Gaussian attributes and camera parameters via graph networks and joint optimization. Furthermore, we incorporate depth cues from optical flow estimation to improve rendering quality. Qualitative results under 6 training views are shown in Fig. 3, where Gesplat not only reconstructs complete geometry but also recovers more accurate and high-frequency details.
Results on Tanks and Temples. To evaluate the performance of Gesplat on large-scale complex scenes, we conduct further comparisons on Tanks and Temples dataset. With a training setting of 6 views, Gesplat achieves the best results across all metrics as shown in Tab. II, demonstrating our robustness on challenging large-scale environments. We also visualize the qualitative comparisons in Fig. 4 to demonstrate the effectiveness of Gesplat.
TABLE III: Ablations on key modules. We evaluate them on the fortress scene.
@@ IV-B Experimental Results and Analysis
Figure 5: Ablation study on the modules of our model.
TABLE IV: Ablations on model parameter settings. We evaluate them on the trex scene.
@@
IV-C Ablation Study
We conduct a series of ablation studies to validate the effectiveness of each module in Gesplat and the rationality of parameter selection. Under a 3 views setting, we select the fortress scene from the LLFF dataset to establish our design choices. Quantitative and qualitative results of the ablation study are presented in Tab. III and Fig. 5, respectively.
Effect of hybrid representation (Hybrid Rep.). As shown in Tab. III, the removal of the hybrid representation leads to a drop in PSNR from 21.14 dB to 18.67 dB. As illustrated in Fig. 5, the visual quality of the rendered image also degrades significantly, with noticeable holes and the blurred texture of the table. These results demonstrate that, compared to the original Gaussian representation, introducing hybrid Gaussian representation based matching priors better preserves geometric information, improves structural completeness, and reduces the risk of overfitting in sparse view scenes.
Effect of graph-guided optimization (Graph Opt.). Although graph-guided optimization is only applied in the last 200 iterations, its removal causes a quantitative decrease of 0.56 dB in PSNR, as shown in Tab. III. From the visual comparison in Fig. 5, we observe that the edges of the floor become blurred and the details of the fortress are oversmoothed without graph-guided optimization. This indicates that graph-guided refinement effectively refines Gaussian attributes and enhances scene details.
Effect of flow-based depth regularization (Depth Reg.). Without flow-based depth regularization, it results in a slight decrease in PSNR, as reported in Tab. III. Compared to our full model, the absence of this module also blurs the edges of the fortress and fails to accurately reconstruct the gaps in the floor, as shown in Fig. 5. These results underscore the importance of this strategy.
Effect of model parameter settings. We further investigate the effectiveness of parameter selection using the trex scene from the LLFF dataset. The parameters under study include the weight of the depth loss $\lambda_{depth}$ and the regularization parameters for Gaussian attribute updates in the graph network: $\lambda_{z}$, $\lambda_{s}$, $\lambda_{r}$, $\lambda_{c}$, and $\lambda_{\alpha}$. The results are summarized in Tab. IV.The value of $\lambda_{depth}$ influences the geometric continuity of the scene. We find that $\lambda_{depth}=0.1$ yields the best performance. Increasing $\lambda_{depth}$ leads to degradation across all metrics, indicating that excessive depth regularization over-smoothens texture details and reduces reconstruction accuracy. Since we expect the graph network to refine scene details through Gaussian attribute updates, we place greater emphasis on SSIM and LPIPS during evaluation. Based on the results in Tab. IV, we set $\lambda_{z}=0.1$, $\lambda_{s}=0.1$, $\lambda_{r}=0.05$, $\lambda_{c}=0.01$, and $\lambda_{\alpha}=1.0$.
V Conclusion
In this paper, we introduce Gesplat to address the challenge of sparse-view novel view synthesis from unposed images. Our framework first employs VGGT to generate an initial dense point cloud and camera pose estimates. To enhance structural consistency, we introduce a hybrid Gaussian representation optimized with matching priors, enforcing geometric coherence through multi-view structural and rendering constraints. We further incorporate flow-based depth regularization to improve rendering accuracy and a graph-guided optimization module to refine Gaussian attributes for detailed scene recovery. Finally, we jointly optimize camera parameters and Gaussian representations. Extensive comparisons with other pose-free methods show that our approach achieves state-of-the-art performance on both forward-facing and large-scale scenes. A limitation remains when handling largely non-overlapping views, where matching priors become unreliable. Future work will explore more robust geometric constraints for extreme sparsity scenarios.

### v2

@@ 3.4.2 Flow-estimated Depth
where $I_{i}$ and $I_{j}$ is a matching image pair. Given a point $p_{i}$ in image $I_{i}$, the predicted corresponding point $\widehat{p}_{j}$ in image $I_{j}$ is estimated as:
@@
Due to estimation errors, $\widehat{p}_{j}$ may not accurately lies on the epipolar line $l_{p_{i}}$. Therefore, we select the perpendicular foot $\overline{p}_{j}$ as the approximate estimation of $\widehat{p}_{j}$ :
@@
where $(\widehat{x}_{j},\widehat{y}_{j})$ is the coordinate of point $\widehat{p}_{j}$, and $\{a,b,c\}$ are the parameters of the epipolar line $l_{p_{i}}$. Then the depth value of the point $p_{i}$ is calculated as:
@@
3.4.3 Cross-view Depth Blending
Given image set $\{I_{i}\mid i=1,2,\cdots,N\}$, for a point $p_{i}$ in $I_{i}$, we compute the predicted matching points $\{\overline{p}_{j}\mid j\neq i\}$ and the corresponding depths in other N1N-1 views. Although $\overline{p}_{j}$ lies on the epipolar line, the distance error that $\overline{p}_{j}$ deviates from the true correspondence $\widehat{p}_{j}$ still persists. To handle the variance of depth estimation across views, we introduce a cross-view depth blending strategy. Depth is represented as the distance information from a pixel point to the camera. Given $O_{i}$ as the camera center of view $i$, we define a reference distance $dis_{ref}$ as the distance from the 3D point $\overline{P}_{j}$ to $O_{i}$. And the projection distance $dis_{pro}$ in view $j$ is the distance from the projected point $\overline{p}_{j}$ to the epipole $o_{ij}$. The distance between $\overline{p}_{j}$ and the true matching point $p_{j}$ is denoted as $\Delta dis_{pro}$, corresponding to a 3D distance $\Delta dis_{ref}$. A smaller $\Delta dis_{ref}$ indicates a more accurate depth estimation. Through the fundamental theorem of calculus, the transformation between $\Delta dis_{ref}$ and $\Delta dis_{pro}$ is computed as:
@@ 3.4.3 Cross-view Depth Blending
where $dis_{ref}^{\prime}(dis_{pro})$ is the gradient of $dis_{ref}(dis_{pro})$, indicating the sensitivity of depth to the changes of $dis_{pro}$. A smaller $dis_{ref}^{\prime}(dis_{pro})$ implies higher confidence in the estimated depth. Using geometric relations, the formulation of $dis_{ref}^{\prime}(dis_{pro})$ is expressed as:
$dis_{ref}^{\prime}(dis_{pro})=\frac{t\sin\beta\sin^{2}(\alpha+\theta)}{m\sin\theta\sin^{2}(\alpha+\beta)},$ (22)
@@
3.4.4 Depth Regularization
Given the rendered depth map $D$ and the flow-estimated depth map $\widehat{D}$, the depth loss is computed using loss function L1:
@@ 3.4.4 Depth Regularization
3.5 Joint Optimization
To jointly optimize noisy point clouds and inaccurate camera poses, we introduce joint optimization, which updates both Gaussians and camera parameters via image loss. Given a Gaussian set $G$ and camera parameters $T$, we minimize the difference between the ground truth image and the rendered image using gradient descent:
@@ 3.5 Joint Optimization
During testing, since predicted camera poses for novel views may still be noisy, we use the trained Gaussian model to refine test camera poses via Eq.25 for 500 iterations per-image for more accurate evaluation.
3.6 Training Objectives
3.6.1 Loss Function
The total training loss is formulated as :
@@ 3.6.1 Loss Function
3.6.2 Training Details
We set $\lambda_{gp}=1.0$, $\lambda_{depth}=0$, $\lambda_{rg}=0$ for the first 2000 iterations, and then adjust to $\lambda_{depth}=0.1$, $\lambda_{rg}=0.3$ to avoid sub-optimization during early stage of training . Our model is built upon InstantSplat [16] and trained for 5000 iterations. The learning rate for the learnable depth $z$ starts at 0.1 and decays to $1.6\times 10^{-6}$. To mitigate initial errors and computational cost from the graph network, we only use it for Gaussian optimization in the final 200 iterations.
4 Experiments
4.1 Experimental Settings
4.1.1 Datasets
Under sparse-view settings, we evaluate Gesplat using LLFF and Tanks and Temples datasets. LLFF provides a collection of real-world scenes with forward-facing camera trajectories, while Tanks and Temples is a large-scale dataset containing both indoor and outdoor scenes with complex environments. Both datasets consist of eight scenes each. Following InstantSplat [16], we randomly sample 24 images from each scene, with 12 images held out as test views (including the first and last images). For the remaining 12 images, we select 3 to 9 images for training under sparse-view conditions.
4.1.2 Baselines &\& Metrics
We compare Gesplat with other pose-free approaches, including Nope-NeRF [14], CF-3DGS [15], and InstantSplat [16]. Among them, Nope-NeRF and CF-3DGS leverage monocular depth estimation and ground-truth camera intrinsics, while InstantSplat relies on DUSt3R [17] to generate dense point clouds. All of the baselines are trained using official code with default settings. We evaluate the performance on novel view synthesis task using common metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [50], and Learned Perceptual Image Patch Similarity (LPIPS) [55].
4.1.3 Implementation Details
Our implementation is built on the PyTorch framework, and all experiments are conducted on one NVIDIA RTX 3090 GPU. For Nope-NeRF and CF-3DGS, we selete input images at the same resolution as original datasets. Using DUSt3R, InstantSplat predicts multi-view stereo depth maps at a resolution of 512. Our method trains with a resolution of 518 × 392 using VGGT [18].
Table 1: Quantitative comparison on LLFF  (3,6,9 views). The best, second best, and third best entries are marked in red, orange, and yellow, respectively.
@@ 4.1.3 Implementation Details
Fig 3: Qualitative comparison between Gesplat and various pose-free methods on LLFF datasets(6 views).The reconstruction of our method is more accurate and exhibits finer details compared with other competitors.
4.2 Experimental Results and Analysis
4.2.1 Results on LLFF
Quantitative comparisons are summarized in Tab. 1. Gesplat achieves the best performance across PSNR, SSIM, and LPIPS compared to other pose-free methods. Nope-NeRF utilizes Multilayer Perceptions (MLPs) to prioritize learning the global scene structure, which effectively suppresses floating artifacts but may produce images suffering from over-blur and require a long training time. CF-3DGS builds upon 3DGS with local and global optimization, enabling more accurate geometric reconstruction and real-time rendering, yet its complex optimization process may lead to loss of fine details. As a feed-forward model, InstantSplat uses DUSt3R to estimate dense point clouds and camera poses, mitigating the information scarcity of 3DGS in sparse views and improving geometric consistency. However, it may occasionally generate inexistent details, reducing the reconstruction quality. By leveraging VGGT for dense 3D geometry recovery, our method uses a hybrid Gaussian representation to learn scene geometry and refines Gaussian attributes and camera parameters via a graph network and joint optimization. Furthermore, we incorporate depth cues from optical flow estimation to improve rendering quality. Qualitative results under 6 training views are shown in Fig. 3, where Gesplat not only reconstructs complete geometry but also recovers more accurate and high-frequency details.
4.2.2 Results on Tanks and Temples
To evaluate the performance of Gesplat on large-scale complex scenes, we conduct further comparisons on Tanks and Temples dataset. With a training setting of 6 views, Gesplat achieves the best results across all metrics as shown in Tab. 2, demonstrating our robustness on challenging large-scale environments. We also visualize the qualitative comparisons in Fig. 4 to demonstrate the effectiveness of Gesplat.
Table 2: Quantitative comparison on Tank and Temples. All models are trained with six input views.The best result is highlighted in bold.
@@ 4.2.2 Results on Tanks and Temples
Fig 4: Qualitative comparison on TNT(3,6,9 views).The reconstruction of our method is more accurate and exhibits finer details.
4.3 Ablation Study
We conduct a series of ablation studies to validate the effectiveness of each module in Gesplat and the rationality of parameter selection. Under a 3 views setting, we select the fortress scene from the LLFF dataset to establish our design choices. Quantitative and qualitative results of the ablation study are presented in Tab. 3 and Fig. 5, respectively.
Table 3: Ablations on the key modules. We evaluate them on the fortress scene of LLFF.
@@ 4.3 Ablation Study
4.3.1 Effect of hybrid representation (Hybrid Rep.)
As shown in Tab. 3, the removal of the hybrid representation leads to a drop in PSNR from 21.14 dB to 18.67 dB. As illustrated in Fig. 5, the visual quality of the rendered image also degrades significantly, with noticeable holes and blurred texture of the table. These results demonstrate that, compared to the original Gaussian representation, introducing hybrid Gaussian representation based matching priors better preserves geometric information, improves structural completeness, and reduces the risk of overfitting in sparse view scenes.
4.3.2 Effect of graph-guided optimization (Graph Opt.)
Although graph-guided optimization is only applied in the last 200 iterations, its removal causes a quantitative decrease of 0.56 dB in PSNR, as shown in Tab. 3. From the visual comparison in Fig. 5, we observe that the edges of the floor become blurred and the details of the fortress are over-smoothed without graph-guided optimization. This indicates that graph-guided optimization effectively refines Gaussian attributes and enhances scene details.
4.3.3 Effect of flow-based depth regularization (Depth Reg.)
Without flow-based depth regularization, it results in a slight decrease in PSNR, as reported in Tab. 3. Compared to our full model, the absence of this module also blurs the edges of the fortress and fails to accurately reconstruct the gaps in the flooring, as shown in Fig. 5. These results underscore the importance of this strategy.
Fig 5: Ablation study on the key modules of our model.
4.3.4 Effect of model parameter settings
We further investigate the effectiveness of parameter selection using the trex scene from the LLFF dataset. The parameters under study include the weight of the depth loss $\lambda_{depth}$ and the regularization parameters for Gaussian attribute updates in the graph network: $\lambda_{z}$, $\lambda_{s}$, $\lambda_{r}$, $\lambda_{c}$, and $\lambda_{\alpha}$. The results are summarized in Tab. 4.The value of $\lambda_{depth}$ influences the geometric continuity of the scene. We find that $\lambda_{depth}=0.1$ yields the best performance. Increasing $\lambda_{depth}$ leads to degradation across all metrics, indicating that excessive depth regularization over-smooths texture details and reduces reconstruction accuracy. Since we expect the graph network to refine scene details through Gaussian attribute updating, we place greater emphasis on SSIM and LPIPS during evaluation. Based on the results in Tab. 4, we set $\lambda_{z}=0.1$, $\lambda_{s}=0.1$, $\lambda_{r}=0.05$, $\lambda_{c}=0.01$, and $\lambda_{\alpha}=1.0$.
Table 4: Ablations on model parameter settings. We evaluate them on the trex scene of LLFF.
@@ 4.3.4 Effect of model parameter settings
5 Conclusion
In this paper, we introduce Gesplat to address the challenge of sparse-view novel view synthesis from unposed images. Our framework first employs VGGT to generate initial dense point clouds and camera pose estimations. To enhance structural consistency, we introduce a hybrid Gaussian representation optimized with matching priors, enforcing geometric coherence through multi-view structural and rendering constraints. We further incorporate flow-based depth regularization to improve rendering accuracy and a graph-guided optimization module to refine Gaussian attributes for detailed scene recovery. Finally, we jointly optimize camera parameters and Gaussian representations. Extensive comparisons with other pose-free methods show that our approach achieves state-of-the-art performance on both forward-facing and large-scale scenes. A limitation remains when handling largely non-overlapping views, where matching priors become unreliable. Future work will explore more robust geometric constraints for extreme sparsity scenarios.